# Extensive entanglement between coupled Tomonaga-Luttinger liquids in and out of equilibrium

### v1

@@ 3 Results: Entanglement and mutual information at finite temperatures
In Fig. 3a we probe the behavior of mutual information $I(A:B)$ in the vicinity of the threshold temperature $T^{*}$. In contrast to decaying entanglement, $I(A:B)$ stays approximately constant. Meanwhile, the magnitude of $I(A:B)$ is comparable to $E_{\mathcal{N}}$ below $T^{*}$, meaning that below the threshold, a significant portion of the correlation consists of quantum entanglement. As one approaches $T^{*}$, the entanglement must get transformed into classical correlations to keep the mutual information constant.
The approximate independence of mutual information with respect to temperature persists far above $T^{*}$ as shown in Fig. 3b. Intuitively, this can be explained if the role of temperature is only to rescale the magnitude of the local fluctuations in $A$ and $B$ without affecting their shared information. Instead, the shared information is governed by the tunneling strength, as supported by $I(A:B)$ monotonically increasing with stronger tunneling. Finally, for the same reason as the entanglement, the mutual information also scales linearly with the subsystem’s length $I(A:B)\propto L$ (Fig. 3c). But unlike entanglement, our analytical and numerical results show that the mutual information does not significantly scale with mean density.
So far, the correlation between the subsystems $A$ and $B$ is induced by tunnel-coupling at the Hamiltonian level and investigated in thermal equilibrium. In the context of experiments in Refs. [17, 11, 18, 12, 19], the temperature needed to observe entanglement between the two gases is an order of magnitude lower than what has been experimentally demonstrated. In the next section, we explore another scenario where the threshold temperature to observe entanglement is within experimental reach. The scenario involves pushing the system far away from equilibrium by coherent splitting.
@@ 4 Results: Entanglement and mutual information after coherent splitting
In the coherent splitting scenario, the system is assumed to be initialized in thermal equilibrium of a single well at $t<0$. Then, the single-well potential is coherently split into a double-well at $t=0$ within a timescale much shorter than the interaction timescale ($\Delta t_{\rm split}\ll\xi_{h}/c$ with $c=\sqrt{gn_{\rm 1D}/m}$ being the speed of sound), such that its effect is analogous to an ideal beam splitter. Let $\hat{b}^{0}$ be the phonon annihilation operator before the splitting, and $\hat{b}^{A,B}$ be the phonon annihilation operators after the splitting, respectively. The case of coherent splitting is analogous to an equal beam splitter with one input being $\hat{b}^{0}$, occupied by a thermal state, and the other input being an ancilla $\hat{b}^{\rm anc}$, occupied by a vacuum. At the output (i.e. after the splitting), all the initial thermal energy gets transferred to the symmetric sector due to the beam splitter operation $\hat{b}^{+}=(\hat{b}^{A}+\hat{b}^{B})/\sqrt{2}=\hat{b}^{0}$. On the other hand, the antisymmetric mode $\hat{b}^{-}=(\hat{b}^{A}-\hat{b}^{B})/\sqrt{2}=\hat{b}^{\rm anc}$ has a minimum uncertainty as it is occupied by the vacuum of the ancilla. If we also reasonably assume $\hat{b}^{0}$ and $b^{\rm anc}$ are initially uncorrelated, it follows that $\hat{b}^{+}$ and $\hat{b}^{-}$ are uncorrelated at the output.
Meanwhile, locally, the atoms undergo a random process of going either to the left or right well, giving excess noise to the relative density $\delta\hat{n}^{-}$ while squeezing the relative phase $\hat{\phi}^{-}$. Phenomenologically, the correlation in the antisymmetric sector after the instantaneous splitting is modelled as a squeezed vacuum of the form [17, 63, 64]
@@
while the phase-density correlations vanish (see e.g. Refs [64, 65] for more theoretically systematic derivations). Here, $r$ is the squeezing parameter, introduced to take into account fluctuation due to technical noise from finite-time splitting [66, 67, 32], or squeezing enhancement accessible in recent experiments [36, 37]. Following the beam splitter argument, the symmetric fluctuation is modelled as thermal with a temperature equal to the condensate before the splitting, and uncorrelated with those in the antisymmetric sector. We remark that although there have been many experimental works on relaxation after coherent splitting, the precise control and characterization of the initial state is a subject of ongoing research [68, 22, 65].
@@ 4.1 Initial entanglement and mutual information at $t=0$
From Eq. (27), we see again a competing behavior: squeezing $r$ and the interaction energy $gn_{\rm 1D}$ (appearing through $\varepsilon_{k}^{+}$) generate entanglement, while temperature appearing through $\eta_{k}^{+}$ suppresses it. For squeezing parameters $r\geq 1$, the first term dominates, whereas for $r\ll 1$, the second term may dominate.
Surprisingly, the threshold temperature $T^{*}$ can be expressed in a closed form
$T^{*}=\sup_{0<k\leq k_{\Lambda}}\left\{T_{k}^{*}=\frac{\varepsilon_{k}^{+}}{k_{B}}\left(\ln\sqrt{\frac{F(k,r)}{F(k,r)-2}}\right)^{-1}\right\}\qquad F(k,r)\equiv 1+\frac{r^{2}}{4}\frac{\varepsilon_{k}^{+}}{E_{k}}+\frac{1}{r^{2}}\frac{E_{k}}{\varepsilon_{k}^{+}}.$ (28)
@@
Figure 4: Entanglement immediately after coherent splitting.— (a) Logarithmic negativity $E_{\mathcal{N}}$ as a function of temperature $T$ for $r=1$ and $n_{\rm 1D}=75\;\rm\mu m^{-1}$. All dashed lines indicate the threshold temperature $T^{*}$. For parameters in panel (a) the threshold temperature is calculated to be $T^{*}\approx 25\;\rm nK$. The solid line is calculated with Eq. (27) and the triangles are computed with numerical diagonalization. Entanglement phase diagram as a function of temperature and (b) mean density $n_{\rm 1D}$ (fixed $r=1$), and (c) squeezing parameter $0.1\leq r\leq 1.5$ (fixed $n_{\rm 1D}=75\;\rm\mu m^{-1}$). The obtained range for $T^{*}$ is accessible in the present day experiments [32, 19]. Both $E_{\mathcal{N}}$ and $T^{*}$ behave non-monotonically with $r$. The condensate length is fixed at $L=100\;\rm\mu m$ for all panels. The condensate length is fixed at $L=100\;\rm\mu m$, while further numerics show that logarithmic negativity still scales linearly with length $E_{\mathcal{N}}\propto L$, as long as the temperature is sufficiently far away from the threshold temperature.
In Fig. 4, we plot $E_{\mathcal{N}}$ and $T^{*}$ for realistic experimental parameters. We first focus on fixed squeezing $r=1$ and mean density $n_{\rm 1D}=75\;\rm\mu m^{-1}$ (Fig. 4a) and calculate the threshold temperature to be $T^{*}\approx 25\;\rm nK$. The threshold temperature can be raised further by increasing interaction energy $gn_{\rm 1D}$ (Fig. 4b) or by increasing squeezing parameter $r$ (Fig. 4c). For realistic range of experimental parameters, $T^{*}$ is found to be between $20\sim 40\;\rm nK$, which is readily accessible in the present day experiments [19, 32].
Moreover, we find that both $E_{\mathcal{N}}$ and $T^{*}$ behave non-monotonically with the squeezing parameter $r$ (Fig. 4c). The system exhibits entanglement for both large ($r>1$) and small ($r<0.2$) squeezing, but the entanglement and the regime of entanglement are strongly suppressed at intermediate squeezing values ($0.2<r<1$). This can be understood from the initial correlation in Eq. (26). As the source of the entanglement is the initial squeezing in the relative phase, setting $r>1$ ($r<1$) would enhance (suppress) the squeezing and thus increase (decrease) the entanglement. But crucially, when $r$ is very small (e.g. $r<0.2$), we are in the opposite regime wherein the density fluctuation is squeezed and so entanglement re-emerge through the second term in Eq. (27).
@@
$I(A:B)=$ $2\sum_{k>0}^{k_{\Lambda}}\left[\left(\lambda_{k}+\frac{1}{2}\right)\log\left(\lambda_{k}+\frac{1}{2}\right)-\left(\lambda_{k}-\frac{1}{2}\right)\log\left(\lambda_{k}-\frac{1}{2}\right)\right]$
@@
$\lambda_{k}=\frac{1}{4}\sqrt{\left(1+2\eta_{k}^{+}+\mathcal{C}_{k}\right)\left(1+2\eta_{k}^{+}+\mathcal{C}_{k}^{-1}\right)}$ (30)
and $\mathcal{C}_{k}=(r^{2}/2)(\varepsilon_{k}^{+}/E_{k})$. Note that the $\lambda_{k}$ above has a similar form as that of Eq. (25) since in both scenarios the covariance matrices have the same structure, differing only in the matrix elements in the antisymmetric sector.
Figure 5: Mutual information immediately after coherent splitting.— (a) The increase of mutual information $I(A:B)$ with increasing temperature $T$ for fixed squeezing $r=1$ and mean density $n_{\rm 1D}=75\;\rm\mu m^{-1}$, in contrast with decaying $E_{\mathcal{N}}$ in Fig. 4a. The solid line is calculated with Eq. (29) and the triangles are computed with numerical diagonalization. (b) The mutual information $I(A:B)$ varies slowly with mean density $n_{\rm 1D}$. Here the squeezing is fixed at $r=1$. (c) Similar to entanglement, $I(A:B)$ also changes non-monotonically with squeezing parameter $r$ ($n_{\rm 1D}=75\;\rm\mu m^{-1}$). In all panels, the dashed lines represent the threshold temperature $T^{*}$, same as in Fig. 4. The condensate length is fixed at $L=100\;\rm\mu m$, while further numerics show that mutual information still scales linearly with length $I(A:B)\propto L$.
In Fig. 5a, we plot $I(A:B)$ as a function of temperature $T$ for fixed squeezing $r=1$ and fixed mean density $n_{\rm 1D}=75\;\rm\mu m^{-1}$. Contrary to the decay of entanglement (Fig. 4a), mutual information grows with temperature irrespective of the threshold. We interpret this as the system being driven further from equilibrium with increasing temperature as the difference between the symmetric and antisymmetric modes becomes more pronounced. This results in a build-up of classical correlations that must not only be enough to compensate for the decaying quantum correlations, but must also add to the overall correlations. This trend does not significantly change as we vary the mean density (Fig. 5b). Similar to entanglement, $I(A:B)$ is also found to behave non-monotonically with squeezing parameter. In Fig. 5c, we observe that $I(A:B)$ is strongly suppressed at low temperatures ($T<20\;\rm nK$) and intermediate squeezing regime ($0.2<r<1$) where entanglement is also suppressed (Fig. 4c).
@@ 4.2 Relaxation constrained by mutual information conservation
In experiments [17, 33], the double-well barrier after coherent splitting is set to be high enough such that no tunneling is allowed between the two gases $A$ and $B$. Consequently, the system’s dynamics follows that of decoupled TLL [Eq. (4) with $J=0$], i.e. $\hat{H}=\hat{H}^{+}_{\rm TLL}+\hat{H}^{-}_{\rm TLL}$. Interestingly, the relative phase fluctuation $\hat{\phi}^{-}(z)$ is experimentally observed to relax into a thermal-like metastable state with an effective inverse temperature $\beta^{-}=(gn_{\rm 1D}/2)^{-1}$, independent of the system’s initial temperature before the splitting [17, 34, 35]. The temperature in the relative sector is inferred from the measured relative phase correlation functions and its full distribution functions [17, 35, 71, 15], whose behavior is found to be consistent with thermal ensemble after relaxation. This phenomenon is an example of prethermalization, widespread in out-of-equilibrium integrable systems [72, 73]. The prethermalization in these experiments arises from the multimode dephasing due to coherent rotation between the relative density and the relative phase fluctuations [17, 24, 32].
Here, we argue that local unitary dynamics generated by decoupled TLL cannot bring the coherently split initial state [Eq. (26)] to a Generalized Gibbs ensemble (GGE) [74, 34] of the form
@@
which is sometimes conjectured to be the relaxed prethermalized state in the past experimental and numerical studies [17, 34, 75], at least approximately. The reason is that the Hamiltonian governing the dynamics is not only decoupled in the $\pm$ basis, but it is also decoupled in $A,B$ basis: $\hat{H}=\hat{H}_{\rm TLL}^{A}+\hat{H}_{\rm TLL}^{B}$. Consequently, the unitary dynamics generated by $\hat{H}$ must conserve the mutual information (and in general, also entanglement) between subsystem $A$ and subsystem $B$, i.e., $I(A:B,t)=I(A:B,0)$ for all $t>0$.
@@
where $\overline{\eta}_{k}=\frac{1}{2}(\eta_{k}^{+}+\eta_{k}^{-})$ is the average of mean thermal occupations in both sectors. By comparing Eq. (32) with Eq. (29), one finds that mutual information is not conserved, hence a contradiction.
@@
Indeed, a detailed theoretical analysis on the relaxation of correlation functions, including the density-density and phase-density correlations, starting from coherently split initial state and following TLL dynamics has been done [64]. It was revealed that the relaxed state is not compatible with a thermal equilibrium, despite some correlation functions agreeing with thermal behavior. Our result highlights another physical constraint, the mutual information conservation, in the subtle process of relaxation and prethermalization of isolated 1D quantum gases.
@@
Alternatively, the system could still prethermalize into GGE of the form (31) in experiments, but it does that not by evolving according to the TLL model. The TLL model is only effective and the exact model almost certainly features non-zero interactions, although the time-scale at which interaction becomes relevant is unclear [76, 18, 77, 78]. In the long-time limit, the system is expected to thermalize at phonon-levels into a thermal state with $\beta^{+}=\beta^{-}=\beta$, having zero mutual information $I(A:B)=0$ for decoupled double-well ($J=0$). This process is expected to occur via anharmonic correction to TLL which include three-body phonon interaction [79, 76, 75, 80, 81]. Experimentally, the physics beyond TLL has not been probed due to the lack of access to the relevant observables [80, 77]. This can now potentially be solved by using common phase measurement [38], giving ways to more detailed and rigorous tests to the GGE hypothesis.
@@
Introducing phonon-phonon interaction into the effective model can lead to interesting dynamics for the entanglement and mutual information studied here. This is because the evolution become effectively non-unitary as excitations may leak to momentum states above the cutoff and to the non-Gaussian degrees of freedom. In experiments, the spatial resolution of the measurements places a cut-off on the momenta that can be resolved [54, 53]. A universal flow of momentum excitations has previously been argued [82]. This means that the dynamics will appear to ‘move’ certain excitations out of view while also ‘move’ some excitations into view. Moreover, interaction may lead to non-Gaussian correlations, whose presence implies that the entropic quantities evaluated under the assumption of the state being Gaussian may not be accurate. However, as long as the non-Gaussian state does not depart too far from being Gaussian, the entropy is continuous under the assumption of a finite energy [83]. If $\rho$ is an arbitrary state, there exists a unique Gaussian state $\rho_{G}$ with the same second moments $\Gamma$ as $\rho$. The entropy computed using the Gaussian formula for $\rho_{G}$ will overestimate the true entropy of $\rho$, i.e. $S(\rho)\leq S(\rho_{G})$. This is because Gaussian entropy disregards higher-order correlations as if they were unobserved, which for a non-Gaussian state corresponds to a lack of knowledge reflected in the higher entropy value. How exactly these momentum cutoff and non-Gaussian correlations affect entanglement and mutual information during relaxation dynamics is a subject for further research.

### v2

@@ 3 Results: Entanglement and mutual information at finite temperatures
We probe the behavior of mutual information $I(A:B)$ in the vicinity of the threshold temperature $T^{*}$ in Fig. 3a. In contrast to decaying entanglement, $I(A:B)$ stays approximately constant. Meanwhile, the magnitude of $I(A:B)$ is comparable to $E_{\mathcal{N}}$ below $T^{*}$, meaning a significant portion of the correlation consists of quantum entanglement. Thus, as one approaches $T^{*}$, the entanglement must get transformed into classical correlations to keep the mutual information constant.
The approximate independence of mutual information with respect to temperature persists far above $T^{*}$ as shown in Fig. 3b. Intuitively, this can be explained if the role of temperature is only to rescale the magnitude of the local fluctuations in $A$ and $B$ without affecting their shared information. Instead, the shared information is governed by the tunneling strength, as supported by $I(A:B)$ monotonically increasing with stronger tunneling. Finally, for the same reason as the logarithmic negativity, the mutual information also scales linearly with the subsystem’s length $I(A:B)\propto L$ (Fig. 3c). But unlike logarithmic negativity, our analytical and numerical results suggest that the mutual information does not significantly scale with mean density.
So far, the correlation between the subsystems $A$ and $B$ is induced by tunnel-coupling at the Hamiltonian level and investigated in thermal equilibrium. In the context of experiments in Refs. [17, 11, 18, 12, 19], the threshold temperature for observing entanglement between the two gases is an order of magnitude lower than what has been experimentally demonstrated. In the next section, we explore another scenario where the threshold temperature is within experimental reach. The scenario involves pushing the system far away from equilibrium by coherent splitting.
@@ 4 Results: Entanglement and mutual information after coherent splitting
In the coherent splitting scenario, the system is assumed to be initialized in thermal equilibrium of a single well at $t<0$. Then, the single-well potential is coherently split into a double-well at $t=0$ within a timescale much shorter than the interaction timescale ($\Delta t_{\rm split}\ll\xi_{h}/c$ with $c=\sqrt{gn_{\rm 1D}/m}$ being the speed of sound), such that its effect is analogous to spatially extended beam splitters. Let $\hat{b}_{k}^{0}$ be the phonon annihilation operator before the splitting (for momentum $k$), and $\hat{b}^{A,B}_{k}$ be the phonon annihilation operators after the splitting, respectively. The case of coherent splitting is analogous to an equal beam splitter with one input being $\hat{b}^{0}_{k}$, occupied by a thermal state, and the other input being an ancilla $\hat{b}^{\rm anc}_{k}$, occupied by a vacuum. At the output (i.e. after the splitting), all the initial thermal energy gets transferred to the symmetric sector due to the beam splitter operation $\hat{b}^{+}_{k}=(\hat{b}^{A}_{k}+\hat{b}^{B}_{k})/\sqrt{2}=\hat{b}^{0}_{k}$. On the other hand, the antisymmetric modes $\hat{b}^{-}_{k}=(\hat{b}^{A}_{k}-\hat{b}^{B}_{k})/\sqrt{2}=\hat{b}^{\rm anc}_{k}$ have minimum uncertainties as they are occupied by the vacuum of the ancilla. If we also reasonably assume that the input $\hat{b}^{0}_{k}$ and $b^{\rm anc}_{k}$ are initially uncorrelated, it follows that $\hat{b}^{+}_{k}$ and $\hat{b}^{-}_{k}$ are uncorrelated at the output.
Meanwhile, locally, the atoms undergo a random process of going either to the left or right well. This random process leads to a binomial density distribution in each small region (of size $\sim\xi_{h}$), giving noise to the local relative density $\delta\hat{n}^{-}(z)$ that is proportional to the mean density $n_{\rm 1D}$ while reducing fluctuation in the relative phase $\hat{\phi}^{-}(z)$ up to the minimum uncertainty relation. Following this argument, the momentum correlation in the antisymmetric sector after the instantaneous splitting is phenomenologically modeled as [17, 73, 74]
@@
while the phase-density correlations vanish (see e.g. Refs [74, 75] for more theoretically systematic derivations). Here, $r$ is the number squeezing parameter, previously used as one of the key global parameters for probing squeezing in 1D gases (denoted $\xi_{N}$ in Refs. [40, 68, 22, 73]). Meanwhile, the symmetric fluctuation is modeled as thermal with a temperature equal to the condensate before the splitting, and uncorrelated with those in the antisymmetric sector. The case where the system is initialized in the ground state before the splitting can be recovered by setting temperature to zero. We remark that although there have been many experimental works on relaxation after coherent splitting, the precise control and characterization of the initial correlation is a subject of ongoing research [76, 22, 75].
@@ 4.1 Initial entanglement and mutual information at $t=0$
From Eq. (28), we see again a competing behavior: squeezing $r$ and the interaction energy $gn_{\rm 1D}$ (appearing through $\varepsilon_{k}^{+}$) generate entanglement, while temperature appearing through $\eta_{k}^{+}$ suppresses it. The first term dominates for squeezing parameters $r\geq 1$, whereas the second term may dominate for $r\ll 1$.
Surprisingly, the threshold temperature $T^{*}$ can be expressed in a closed form (see B for derivation),
$T^{*}=\sup_{0<k\leq k_{\Lambda}}\left\{T_{k}^{*}=\frac{\varepsilon_{k}^{+}}{k_{B}}\left(\ln\sqrt{\frac{F(k,r)}{F(k,r)-2}}\right)^{-1}\right\}\qquad F(k,r)\coloneqq 1+\frac{r^{2}}{4}\frac{\varepsilon_{k}^{+}}{E_{k}}+\frac{1}{r^{2}}\frac{E_{k}}{\varepsilon_{k}^{+}}.$ (29)
@@
Figure 4: Entanglement immediately after coherent splitting.— (a) Logarithmic negativity $E_{\mathcal{N}}$ as a function of temperature $T$ for $r=1$ and $n_{\rm 1D}=75\;\rm\mu m^{-1}$. All dashed lines indicate the threshold temperature $T^{*}$. For parameters in panel (a) the threshold temperature is calculated to be $T^{*}\approx 25\;\rm nK$. The solid line is calculated with Eq. (28) and the triangles are computed with numerical diagonalization. Entanglement phase diagram as a function of temperature and (b) mean density $n_{\rm 1D}$ (fixed $r=1$), and (c) squeezing parameter $0.1\leq r\leq 1.5$ (fixed $n_{\rm 1D}=75\;\rm\mu m^{-1}$). The obtained range for $T^{*}$ is accessible in the present day experiments [32, 19]. Both $E_{\mathcal{N}}$ and $T^{*}$ behave non-monotonically with $r$. The condensate length is fixed at $L=100\;\rm\mu m$ for all panels, while further numerics show that logarithmic negativity still scales linearly with length $E_{\mathcal{N}}\propto L$, as long as the temperature is sufficiently far away from the threshold temperature.
In Fig. 4, we plot $E_{\mathcal{N}}$ as a function of $T$ for realistic experimental parameters. We first focus on fixed squeezing $r=1$ and mean density $n_{\rm 1D}=75\;\rm\mu m^{-1}$ (Fig. 4a) and calculate the threshold temperature to be $T^{*}\approx 25\;\rm nK$. The threshold temperature can be raised further by increasing interaction energy $gn_{\rm 1D}$ (Fig. 4b) or by increasing squeezing parameter $r$ (Fig. 4c). For realistic range of experimental parameters, $T^{*}$ is found to be between $20\sim 40\;\rm nK$, which is readily accessible in the present day experiments [19, 32].
Moreover, we find that both $E_{\mathcal{N}}$ and $T^{*}$ behave non-monotonically with the squeezing parameter $r$ (Fig. 4c). The system exhibits entanglement for both large ($r>1$) and small ($r<0.2$) squeezing, but the measure and the regime of entanglement are strongly suppressed at intermediate squeezing values ($0.2<r<1$). This can be understood from the initial correlation in Eq. (27). As the source of the entanglement is the strong correlation between the two phase fields (suppression of the relative phase fluctuation), setting $r>1$ ($r<1$) would enhance (suppress) this correlation. But when $r$ is very small (e.g. $r<0.2$), we are in the opposite regime wherein the density fields are strongly correlated (relative density fluctuation is suppressed) and so entanglement re-emerge through the second term in Eq. (28).
@@
$I(A:B)=$ $\sum_{k>0}^{k_{\Lambda}}2\left[\left(\lambda_{k,r}+\frac{1}{2}\right)\log\left(\lambda_{k,r}+\frac{1}{2}\right)-\left(\lambda_{k,r}-\frac{1}{2}\right)\log\left(\lambda_{k,r}-\frac{1}{2}\right)\right]$
@@
$\lambda_{k,r}=\frac{1}{4}\sqrt{\left(1+2\eta_{k}^{+}+\mathcal{C}_{k,r}\right)\left(1+2\eta_{k}^{+}+\mathcal{C}_{k,r}^{-1}\right)}$ (31)
and $\mathcal{C}_{k,r}\coloneqq(r^{2}/2)(\varepsilon_{k}^{+}/E_{k})$. In the case where the system is prepared in the ground state before the splitting, entanglement entropy can be calculated from Eq. (30) and Eq. (31) by setting $\eta_{k}^{+}=0$. Note that the $\lambda_{k,r}$ above has a similar form as that of Eq. (26) since in both scenarios the covariance matrices have the same structure, differing only in the matrix elements in the antisymmetric sector.
Figure 5: Mutual information immediately after coherent splitting.— (a) The increase of mutual information $I(A:B)$ with increasing temperature $T$ for fixed squeezing $r=1$ and mean density $n_{\rm 1D}=75\;\rm\mu m^{-1}$, in contrast with decaying $E_{\mathcal{N}}$ in Fig. 4a. The solid line is calculated with Eq. (30) and the triangles are computed with numerical diagonalization. (b) The mutual information $I(A:B)$ varies slowly with mean density $n_{\rm 1D}$. Here the squeezing is fixed at $r=1$. (c) Similar to entanglement, $I(A:B)$ also changes non-monotonically with squeezing parameter $r$ ($n_{\rm 1D}=75\;\rm\mu m^{-1}$). In all panels, the dashed lines represent the threshold temperature $T^{*}$, same as in Fig. 4, and the condensate length is fixed at $L=100\;\rm\mu m$, while further numerics show that mutual information still scales linearly with length $I(A:B)\propto L$.
In Fig. 5a, we plot the mutual information $I(A:B)$ as a function of temperature $T$ for fixed squeezing $r=1$ and fixed mean density $n_{\rm 1D}=75\;\rm\mu m^{-1}$. Contrary to the decay of entanglement (Fig. 4a), $I(A:B)$ grows with $T$ irrespective of the threshold, the trend we also observe as we vary $n_{\rm 1D}$ (Fig. 5b). We interpret this as the system being driven further from equilibrium with increasing $T$, as the difference between the symmetric and antisymmetric modes after the splitting becomes more pronounced. Thus, the splitting produces both quantum and classical correlations, with the classical correlation dominate as the temperature increases. Similar to entanglement, $I(A:B)$ is also found to behave non-monotonically with the squeezing parameter $r$. In Fig. 5c, we observe that $I(A:B)$ is strongly suppressed at low temperatures ($T<20\;\rm nK$) and intermediate squeezing regime ($0.2<r<1$), where entanglement is also suppressed (Fig. 4c).
@@ 4.2 Relaxation constrained by mutual information conservation
In experiments [17, 33], the double-well barrier after coherent splitting is set to be high enough such that no tunneling is allowed between the two gases $A$ and $B$. Consequently, the system’s dynamics follows that of decoupled TLL [Eq. (4) with $J=0$], i.e. $\hat{H}=\hat{H}^{+}_{\rm TLL}+\hat{H}^{-}_{\rm TLL}$. Interestingly, the relative phase fluctuation $\hat{\phi}^{-}(z)$ is experimentally observed to relax into a thermal-like metastable state with an effective inverse temperature $\beta^{-}=(gn_{\rm 1D}/2)^{-1}$, independent of the system’s initial temperature before the splitting [17, 34, 35]. The temperature in the relative sector is inferred from the measured relative phase correlation functions and its full distribution functions [17, 35, 79, 15], whose behavior is found to be consistent with thermal ensemble after relaxation. This phenomenon is an example of prethermalization, widespread in out-of-equilibrium integrable systems [80, 81]. The prethermalization in these experiments is modeled by multimode dephasing due to coherent rotation between the relative density and the relative phase fluctuations [17, 24, 32].
On the contrary, here we argue that local unitary dynamics generated by decoupled TLL cannot bring the coherently split initial state [Eq. (27)] to a Generalized Gibbs ensemble (GGE) [82, 34] of the form
@@
which is often conjectured to be the relaxed prethermalized state in the past experimental and numerical studies [17, 34, 83]. Here we focus specifically on the GGE of the form Eq. (32) without ruling out other forms of GGE. The state (32) cannot be reached under TLL dynamics because the Hamiltonian not only decouples in the $\pm$ basis but it is also decoupled in the $A,B$ basis: $\hat{H}=\hat{H}_{\rm TLL}^{A}+\hat{H}_{\rm TLL}^{B}$. Consequently, the unitary dynamics generated by $\hat{H}$ must conserve the mutual information (and in general, also entanglement) between subsystem $A$ and subsystem $B$, i.e., $I(A:B,t)=I(A:B,0)$ for all $t>0$.
@@
where $\overline{\eta}_{k}=\frac{1}{2}(\eta_{k}^{+}+\eta_{k}^{-})$ is the average of thermal occupations in both sectors. By comparing Eq. (33) with Eq. (30), one finds that mutual information is not conserved, hence a contradiction.
@@
Since the symmetric sector is assumed to be always thermal, our results show that the antisymmetric sector cannot reach the Gibbs state from the given initial correlation (27) via TLL dynamics. Indeed, the relaxation of correlation functions in the antisymmetric sector, including density-density correlation, after a coherent split has been theoretically analyzed in detail [74]. It was revealed that the correlation functions are not exactly compatible with a thermal equilibrium in the long-time limit. However, numerically, they look very similar that they may practically be indistinguishable in experiments. Our result highlights another physical constraint—the mutual information conservation—which extends beyond the behavior captured by correlation functions in the subtle process of relaxation and prethermalization in isolated 1D quantum gases.
@@
Alternatively, the system could still prethermalize into GGE of the form (32) in experiments, but at least one of the following must be true: the initial correlation is not of the form Eq. (27), or the system does not evolve according to the TLL model. The TLL model is only effective and the exact model almost certainly features non-zero interactions, although the time-scale at which interaction becomes relevant is unclear [84, 18, 85, 86]. In the long-time limit, the system is expected to thermalize at phonon-levels into a thermal state with $\beta^{+}=\beta^{-}=\beta$, having zero mutual information $I(A:B)=0$ for decoupled double-well ($J=0$). This process is expected to occur via anharmonic correction to TLL which include three-body phonon interaction [87, 84, 83, 88, 89].
@@
Introducing phonon-phonon interaction into the effective model can lead to interesting dynamics for the entanglement and mutual information studied here. This is because the evolution become effectively non-unitary as excitations may leak to momentum states above the cutoff and to the non-Gaussian degrees of freedom. In experiments, the spatial resolution of the measurements places a cut-off on the momenta that can be resolved [46, 45]. A universal flow of momentum excitations has previously been argued [90]. This means that the dynamics will appear to ‘move’ certain excitations out of view while also ‘move’ some excitations into view. Moreover, interaction may lead to a build up of non-Gaussian correlations, implying that the entropic quantities evaluated under the assumption of the state being Gaussian may not be accurate. However, as long as the non-Gaussian state does not depart too far from being Gaussian, the entropy is continuous under the assumption of a finite energy [91]. If $\rho$ is an arbitrary state, there exists a unique Gaussian state $\rho_{G}$ with the same second moments $\Gamma$ as $\rho$. The entropy computed using the Gaussian formula for $\rho_{G}$ will overestimate the true entropy of $\rho$, i.e. $S(\rho)\leq S(\rho_{G})$. This is because Gaussian entropy disregards higher-order correlations as if they were unobserved, which for a non-Gaussian state corresponds to a lack of knowledge reflected in the higher entropy value. How exactly these momentum cutoff and non-Gaussian correlations affect entanglement and mutual information during relaxation dynamics is a subject for further research.